Circular No. 8699

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 G1 (McNAUGHT) R. H. McNaught reports his discovery of a comet with a coma diameter

of ∼ 10′′ on CCD images taken with the 0.5-m Uppsala Schmidt telescope in poor, variable seeing conditions (discovery observation below). Followup observations by McNaught on Apr. 6.6–6.7 UT showed a faint, diffuse tail ∼ 8′′ long in p.a. ∼ 300^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Apr. 5.69705  14^{h}31^{m}48.^{s}10  −42^{o}58′12′′.3  18.0

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-G27. It is likely that this comet is of short period.

T = 2006 Oct. 10.798 TT  ω = 17.^{o}146
Ω = 283.217 }2000.0
q = 1.59344 AU  i = 20.334

SUPERNOVAE 2006bm, 2006bn, 2006bo Three additional apparent supernova have been reported from unfil-

tered CCD images: 2006bm by T. Puckett and G. Sostero (0.60-m reflector at Ellijay, GA; cf. IAUC 8693), 2006bn by L. A. G. Monard (cf. IAUC 8697), and 2006bo by T. Boles (cf. IAUC 8698).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006bm Apr. 5.195 11^{h}18^{m}21.^{s}41 +28^{o}12′43′′.5 18.1 4′′.1 W, 10′′.5 N
2006bn Apr. 5.087 20 00 15.75 −38 34 18.3 18.4 42′′ E, 24′′ N
2006bo Apr. 5.148 20 30 41.90 + 9 11 40.8 17.6 12′′ W, 14′′ N

Additional magnitudes for 2006bm: 2005 Sept. 10 UT, [19.4; 2006 Apr. 6.125, 17.9. Additional magnitudes for 2006bn in IC 4926: 2005 Sept. 21.25, [19.0; 2006 Apr. 6.094, 18.4 ± 0.3; 7.111, 18.6 ± 0.2. Nothing is visible at the location of 2006bn on the Digitized Sky Survey (limiting red mag 20.5). Additional magnitudes for 2006bo in UGC 11578: 2005 Oct. 9 and Nov. 28, [19.5 (Boles); 2006 Apr. 7.137, 17.1 (Monard); 7.151, 17.6 (Boles). Monard’s image yields position end figures 41^{s}.95, 40′′.8 for 2006bo. Nothing is visible at the location of 2006bo on Digitized Sky Survey plates from 1990 Aug. 20 (limiting red mag 20.5) and 1998 June 15 (limiting blue mag 21.0).

2006 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green